OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point. m=4, (6,5)

OpenStudy (anonymous):

y-5=4(x-6)

OpenStudy (anonymous):

P/s from your equation, is y=4x-24-5 y=4x-19 or y=4x+19

OpenStudy (anonymous):

y-5=4x-24 which gives y=4x-19